(4z^2+4z-80)/(z^2-11z+28) Simplify
Answer by mananth(16946)

(4z^2+4z-80)/(z^2-11z+28) Simplify
(4z^2+4z-80)
= 4z^2+20z-16z-80
=4z(z+5)-16(z+5)
=(4z-16)(z+5)
=4(z-4)(z+5)

(z^2-11z+28)
=z^2-4z-7z+28
=z(z-4)-7(z-4)
=(z-4)(z-7)
=(z-4)(z-7)
4(z-4)(z+5)/ (z-4)(z-7)
=4(z+5)/(z-7)